OpenStudy (hippoharry):

Which of the following represents a first step in solving the equation (x - 7)^2 = 100? A. x - 7 = 100 B. x - 49 = 10 C. x - 7 = 10 D. (x - 7)^2 = 10

OpenStudy (calculusxy):

Normally in completing the square, what you would do in the next step is find the square root of both sides to get rid of the second power next to (x - 7). \(\sqrt{(x-7)^2} = \sqrt{100}\) Then you would have \(x - 7 = \pm 10\)

OpenStudy (tt0860):

pemdas

OpenStudy (tt0860):

parenthesis

OpenStudy (calculusxy):

@hippoharry Do you understand how this was done?